# Non-Hermitian Hamiltonian Approach for Two-Dimensional Spectroscopy

### v1

@@ Abstract
Two-dimensional spectroscopy (2DS) offers significant advantages in terms of high temporal and frequency resolutions and signal-to-noise ratio. Until now, the response-function (RF) formalism has been the prevalent theoretical description. In this study, we compare the non-Hermitian Hamiltonian (NHH) method with the RF formalism in a three-level system with a constant control field. We obtain the signals from both approaches and compare their population dynamics and 2DS. We propose the quasi-Green function for the NHH method, which allows all possible Liouville paths to be inferred. Although the NHH method overestimates relaxations, it also provides a more comprehensive description. Our results demonstrate that the NHH method is more suitable than the RF formalism for investigating the systems that are either dissipative or complex via the 2DS.
@@
Two-Dimensional Spectroscopy,Non-Hermitian Hamiltonian Method,Response-Function Formalism,
@@ 1 Introduction
Two-dimensional spectroscopy (2DS) is a technique that applies sequential pulses to materials or biological systems and receives output signals rich in structural and dynamic information 1, 2, 3, 4. This technique is implemented with a wide range of pulse wavelengths, from UV to microwave, providing temporal resolutions ranging from milliseconds to attoseconds 5, 6, 7, 8, 9. It is widely employed in physical chemistry, condensed-matter physics, and biophysics for exploring phenomena such as energy transfer 10, 11, 12, 13, molecular vibrations 14, 15, electronic transitions and relaxation 16, 17, 18, and chemical reactions 19, 20, as well as structures like proteins 21, 22. The means to implement 2DS experimentally are diverse 23, 24, 25, 26, 27, 28. The common principle involves the application of pulse sequences with different time intervals to samples, in which the pulses are well separated. The emitted polarization signals are projected onto two independent frequency axes, creating a map known as 2DS.
@@
On the other hand, the most widely-used method for the theoretical description of the 2DS is the response-function (RF) formalism 29, 30, 31. This approach employs the response functions, each corresponding to a Liouville path, to describe the polarization. The signals are viewed as the responses to the perturbations induced by the three applied pulses. However, it may overlook some Liouville paths if the system setup is complex. For example, an additional field is applied to induce transitions in the system. It may become exhausting to draw all the possible paths, although most of them do not contribute significantly. On the other hand, the non-Hermitian Hamiltonian (NHH) method, which is widely-used in investigating PT symmetry 32, 33, 34, photosynthetic light harvesting 35, 36, 37, chemical reaction and avian compass 38, 39, 40, and negative refraction in Möbius molecules 41, may overcome these disadvantages. By omitting the quantum-jump terms, it obtains a reliable solution to the quantum master equation and thus achieves a balance between computational difficulty and comprehensive description 42. Nevertheless, working with non-Hermitian operators still requires careful consideration in interpreting the physical processes underlying the results.
@@
In this article, we compare the NHH method and the RF formalism in a three-level system with a constantly-applied control field. Although Lemmer et al. first applied the NHH method to 2DS 42, they did not focus on the approach itself but rather on its application to study ion Coulomb crystals. In contrast, our work concentrates on the similarities and differences between these two methods in calculating 2DS. We validate their consistency by comparing the 2DS when they follow identical Liouville paths. We discuss their respective advantages and disadvantages, as well as their scopes of application. Particularly, we propose the quasi-Green function for interpreting the 2DS by the NHH method, which significantly simplifies the analytical challenges. Overall, we believe that the NHH method provides a reliable description for 2DS, particularly in complex systems, and can overcome the limitations of the RF formalism in these scenarios.
@@
We first illustrate the RF formalism and the NHH method, in which the quasi-Green function is proposed as a counterpart to the Green function in the RF formalism. We then present their double-sided Feynman diagrams and the analytical expressions of the rephasing (RP) signal. We further provide a comparison of the Green and quasi-Green functions, the population dynamics, the conservation of the population and the 2DS by the two approaches, focusing mainly on their differences in coherent and incoherent behavior. The advantages and disadvantages, as well as the difference and consistency of the two methods are discussed. Finally, we summarize our results and offer a perspective on the application of the NHH method.
@@ 2.1 Response Function Formalism
We consider a three-level system including a ground state $|b\rangle$, an excited state $|e\rangle$, and a metastable state $|c\rangle$, as shown in Fig. 1(a). The control field with Rabi frequency $\Omega_{c}$ is constantly applied to the system. When a probe field is introduced, the electromagnetically-induced transparency (EIT) effect is formed to suppress the dissipation due to quantum interference 43. For 2DS, three incident pulses of the probe field induce the $|b\rangle$leftrightarrow $|e\rangle$ transition with Rabi frequency $\Omega_{p}$, and the signals to be measured are emitted from the system in a specific direction, as shown in Fig. 1(b). Between the energy levels $b$ and $e$, there exist bidirectional relaxations with a downhill rate $\Gamma_{1}$ and an uphill rate $\Gamma_{2}$. Basically, the Hamiltonian of the system without probe and control fields is
@@
$(j=b,e,c)$ where we assume $\hbar=1$ for simplicity. Therefore, the interaction Hamiltonian under the control field is written as
$H_{\rm{int}}=\Omega_{c}|e\rangle\langle c|+{\rm h.c.}$ (2)
where $\Omega_{c}=-\frac{1}{2}\Omega_{ec}e^{-i\nu_{c}t}$ with $\Omega_{ec}$ and $\nu_{c}$ being the Rabi frequency and the frequency of the control field respectively. Generally speaking, the probe pulses are strong but short enough to be considered perturbations to the system.
Figure 1: (a) Schematic diagram of the three-level system. The energy levels $b$, $e$ and $c$ are the ground, the excited and the metastable states, respectively. The probe pulses denoted as the light-green line are applied to induce the $|b\rangle$leftrightarrow $|e\rangle$ transition with Rabi frequency $\Omega_{p}$ $(p=a,b,c)$, and the control field denoted as the dark-green line is resonantly coupled to the $|e\rangle$leftrightarrow $|c\rangle$ transition with Rabi frequency $\Omega_{ec}$. (b) Three probe pulses and the signal. (c) Sequence of pulses and signal, in which $t_{1}$, $t_{2}$ and $t_{3}$ are respectively the time intervals between two adjacent pulses. The dashed line represents the emitted signal.
The 2DS are plotted using the third-order response function after double Fourier transformation. This function is proportional to the intensity of third-order polarization $\langle\mu\rho^{(3)}(t)\rangle$ and is expressed as
@@
$\langle\mu\left(t_{3}+t_{2}+t_{1}\right)\left[\mu\left(t_{2}+t_{1}\right),\left[\mu\left(t_{1}\right),\left[\mu\left(0\right),\rho\left(0\right)\right]\right]\right]\rangle,$
where $t_{1}$, $t_{2}$ and $t_{3}$ are the time intervals between two adjacent pulses, as shown in Fig. 1(c). $\rho(0)$ represents the initial equilibrium state under the Hamiltonian $H_{0}$, while $\mu(s)=e^{iH_{0}s}\mu e^{-iH_{0}s}$ is the dipole operator in the interaction picture. By expanding the commutators, the third-order polarization is divided into eight terms with
$\begin{split}R_{1}=&\langle\mu(t_{3}+t_{2}+t_{1})\mu(0)\rho(0)\mu(t_{1})\mu(t_{2}+t_{1})\rangle,\\ R_{2}=&\langle\mu(t_{3}+t_{2}+t_{1})\mu(t_{1})\rho(0)\mu(0)\mu(t_{2}+t_{1})\rangle,\\ R_{3}=&\langle\mu(t_{3}+t_{2}+t_{1})\mu(t_{2}+t_{1})\rho(0)\mu(0)\mu(t_{1})\rangle,\\ R_{4}=&\langle\mu(t_{3}+t_{2}+t_{1})\mu(t_{2}+t_{1})\mu(t_{1})\mu(0)\rho(0)\rangle.\end{split}$ (4)
and their hermitian conjugates. The incidence of three probe pulses from different directions allows for selective observation of certain combinations of the eight terms mentioned above. For example, the RP signal is emitted in the direction $\vec{k}_{s}=-\vec{k}_{a}+\vec{k}_{b}+\vec{k}_{c}$ as
@@
$R_{n}$ $(n=1,2,3,4)$ and their hermitian conjugates, can be obtained by solving the quantum master equation. This equation describes quantum dynamics of the open system and is written as
@@
where $A_{ij}=|i\rangle\langle j|$ is the quantum jump operator describing the process of jumping from $|j\rangle$ to $|i\rangle$, $\gamma_{j}^{\left(0\right)}$ is the pure-dephasing noise for each state, and $\mathfrak{L}\left(A_{ij}\right)\rho=\frac{1}{2}\left\{A_{ij}^{\dagger}A_{ij},\rho\right\}-A_{ij}\rho A_{ij}^{\dagger}$ with $\left\{A_{ij}^{\dagger}A_{ij},\rho\right\}=A_{ij}^{\dagger}A_{ij}\rho+\rho A_{ij}^{\dagger}A_{ij}$ being the anti-commutator. Assuming that the control field is in resonance with the $|e\rangle\leftrightarrow|c\rangle$ transition, the time evolution of the elements of the density matrix in the interaction picture are as follows
$\begin{split}\dot{\rho}_{bb}=&\Gamma_{1}\rho_{ee}-\Gamma_{2}\rho_{bb},\\ \dot{\rho}_{ee}=&\frac{i}{2}\Omega_{ec}\left(\rho_{ce}-\rho_{ec}\right)-\Gamma_{1}\rho_{ee}+\Gamma_{2}\rho_{bb},\\ \dot{\rho}_{cc}=&-\frac{i}{2}\Omega_{ec}\left(\rho_{ce}-\rho_{ec}\right),\\ \dot{\rho}_{eb}=&\frac{i}{2}\Omega_{ec}\rho_{cb}-\gamma_{eb}\rho_{eb},\\ \dot{\rho}_{be}=&-\frac{i}{2}\Omega_{ec}\rho_{bc}-\gamma_{eb}\rho_{be},\\ \dot{\rho}_{ec}=&\frac{i}{2}\Omega_{ec}\left(\rho_{cc}-\rho_{ee}\right)-\gamma_{ec}^{+}\rho_{ec},\\ \dot{\rho}_{ce}=&-\frac{i}{2}\Omega_{ec}\left(\rho_{cc}-\rho_{ee}\right)-\gamma_{ec}^{+}\rho_{ce},\\ \dot{\rho}_{cb}=&\frac{i}{2}\Omega_{ec}\rho_{eb}-\gamma_{bc}\rho_{cb},\\ \dot{\rho}_{bc}=&-\frac{i}{2}\Omega_{ec}\rho_{be}-\gamma_{bc}\rho_{bc},\end{split}$ (7)
@@
Using the Green functions to further describe the time evolution of the density matrix elements, the RP signal is written in a more specific form 44
@@
where the Green function $\mathcal{G}_{ijkl}(t)=\rho_{ij}(t)$ for the initial condition $\rho_{mn}(0)=\delta_{mk}\delta_{nl}$. The double-sided Feynman diagrams for each term in Eq. (11) are shown in Fig. 2, with (a), (b), (c) and (d) corresponding to four terms in turn. The evolutions during $t_{2}$ with the initial and final states are explicitly demonstrated in the diagrams. The first and the last terms of Eq. (11) respectively correspond to $R_{2}$ and $R_{3}$ in Eq. (4). The other two terms, however, do not correspond to $R_{n}$ because they are induced by the population transfer during $t_{2}$. The evolution of the non-diagonal elements in $t_{1}$ and $t_{3}$ are described by the Green functions as
@@
while the evolution in $t_{2}$ is obtained by assuming two possible initial states after the first two pulses. In the situation where the system is at $|e\rangle$ at the beginning of the population time, the Green functions for the system to remain at $|e\rangle$ or to evolve to $|b\rangle$ are respectively written as
@@
Eventually, by taking the real part of the double Fourier transformed response function, the RP signal and thus the 2DS are obtained.
@@
We apply the NHH method to the same system in Fig. 1(a). Starting directly from the master Eq. (6), we drop the quantum-jump term and approximately obtain
@@
where the non-Hermitian Hamiltonian reads
@@
The NHH method is valid as long as the dissipation and dephasing rates $\Gamma_{j}$ and $\gamma_{j}^{(0)}$ are much smaller than the typical frequency of the Hamiltonian.
By the NHH method, the final state of the system in the interaction picture is written as
$|\psi(t_{3},t_{2},t_{1})\rangle=U^{I}_{f}(t_{1}+t_{2}+t_{3}+\delta t_{a}+\delta t_{b}+\delta t_{c},$
@@
×$\times U^{I}_{p}\left(\delta t_{a},0\right)|\psi\left(0\right)\rangle,$ (22)
where $U^{I}_{p}$ ($U^{I}_{f}$) represents the evolution operator with (without) the probe pulses included in the interaction picture. We consider that the probe pulses are so strong and their durations are so short that both the dissipation and dephasing during the excitations can be neglected. In the interaction picture, the Hamiltonians during probe pulses are explicitly written as
$H^{I}_{p}=$ $-\frac{1}{2}\Omega_{be}e^{-i\vec{k}_{p}\cdot\vec{r}}|b\rangle\langle e|+{\rm h.c.},$ (23)
while the Hamiltonians during the free evolutions read
@@
Hereinto, we approximate the probe pulses as square pulses. The time evolutions under $H^{I}_{f}$ for initial states $|b\rangle$, $|e\rangle$ and $|c\rangle$ are respectively written as
$U_{f}^{I}(t_{c})|b\rangle$ $=C_{bb}(t_{c})|b\rangle,$ (25)
$U_{f}^{I}(t_{c})|e\rangle$ $=C_{ee}(t_{c})|e\rangle+C_{ce}(t_{c})|c\rangle,$ (26)
$U_{f}^{I}(t_{c})|c\rangle$ $=C_{ec}(t_{c})|e\rangle+C_{cc}(t_{c})|c\rangle,$ (27)
where $t_{c}~{}(c=1,2,3)$ are time intervals, and
@@
Meanwhile, the time evolutions under $H^{I}_{p}$ for three initial states are respectively written as
$U_{p}^{I}(\delta t_{p})|b\rangle=$ $N_{p}\left(|b\rangle+\beta_{p}e^{i\vec{k_{p}}\vec{r}}|e\rangle\right),$ (36)
$U_{p}^{I}(\delta t_{p})|e\rangle=$ $N_{p}\left(\beta_{p}e^{-i\vec{k_{p}}\vec{r}}|b\rangle+|e\rangle\right),$ (37)
$U_{p}^{I}(\delta t_{p})|c\rangle=$ $|c\rangle,$ (38)

### v2

# Non-Hermitian Hamiltonian Approach for Two-Dimensional Coherent Spectra of Driven Systems

@@ Abstract
Two-dimensional coherent spectroscopy (2DCS) offers significant advantages in terms of high temporal and frequency resolutions and signal-to-noise ratio. Until now, the response-function (RF) formalism has been the prevalent theoretical description. In this study, we compare the non-Hermitian Hamiltonian (NHH) method with the RF formalism in a three-level system with a constant control field. We obtain the signals from both approaches and compare their population dynamics and 2DCS. We propose the quasi-Green functions for the NHH method, which allows all dominant Liouville paths to be inferred. We further simulate the 2DCS of Rh(CO)2C5H7O2 (RDC) dissolved in hexane with the NHH method, which is in good agreement with the previous experiments. Although the NHH method overestimates relaxations, it provides all important paths by analytical solutions, which are different from the four paths used in the RF formalism. Our results demonstrate that the NHH method is more suitable than the RF formalism for investigating the systems including relaxation and control fields via the 2DCS.
@@
Two-Dimensional Coherent Spectroscopy,Non-Hermitian Hamiltonian Method,Response-Function Formalism,
@@ 1 Introduction
Two-dimensional coherent spectroscopy (2DCS) is a technique that applies sequential pulses to materials or biological systems and receives output signals rich in structural and dynamic information 1, 2, 3, 4, 5. In 1998, Jonas et al. first experimentally demonstrated two-dimensional electronic spectroscopy 6. Subsequently, Asplund et al. demonstrated two-dimensional infrared spectroscopy 7. This technology has rapidly advanced in both experimental methods and theoretical approaches since it was first proposed. Nowadays, this technique is implemented with a wide range of pulse wavelengths, from UV to microwave 6, 8, 9, 10, 11, 12, 13, 14, 15, providing temporal resolutions ranging from milliseconds to attoseconds 9, 16, 17, 18, 19. It is widely employed in physical chemistry, condensed-matter physics, and biophysics for exploring phenomena such as energy transfer 20, 21, 22, 23, 24, molecular vibrations 25, 26, electronic transitions and relaxations 27, 28, 29, chemical reactions 30, 31, as well as structures like proteins 32, 33, 34. The means to implement 2DCS experimentally are diverse 35, 36, 37, 38, 39, 40. The common principle involves the application of pulse sequences with different time intervals to samples, in which the pulses are well separated. The emitted polarization signals are projected onto two independent frequency axes, creating a map known as 2DCS.
@@
On the other hand, one of the widely-used methods for the theoretical description of the 2DCS is the response-function (RF) formalism 41, 42, 43, 44. This approach employs the double-sided Feynman diagrams, each corresponding to a Liouville path, to describe the polarization. The signals are viewed as the responses to the perturbations induced by the three applied pulses. However, the number of diagrams to be drawn and calculated may reach as many as hundreds if the system involves more external fields and energy levels. In such cases, the RF formalism may lead to the omission of some important diagrams. For example, an additional field is applied to induce transitions in the system. It may become exhausting to draw all possible paths to obtain the analytical solution, although most of them do not contribute significantly. On the other hand, the non-Hermitian Hamiltonian (NHH) method, which is widely-used in investigating PT symmetry 45, 46, 47, photosynthetic light harvesting 48, 49, 50, chemical reaction and avian compass 51, 52, 53, and negative refraction in Möbius molecules 54, may overcome these disadvantages. By omitting the quantum-jump terms, it obtains a reliable solution to the quantum master equation and thus achieves a balance between computational difficulty and analytical-solution-based description 55. Nevertheless, working with non-Hermitian operators still requires careful consideration in interpreting the physical processes underlying the results. Despite the RF formalism and the NHH method, other approaches with different advantages for simulating 2DCS include the hierarchical equations of motion approach 56, the nonperturbative response function formalism 57, the numerical integration of the Schrödinger equation scheme 58, 59, the non-linear exciton equations method 2, the non-linear exciton propagation method 60 and the full cumulant expansion formalism 61.
@@
In this article, we compare the NHH method and the RF formalism in a three-level system with a constantly-applied control field. It is a typical driven system, characterized by additional electromagnetic fields beyond the pulse sequence used in 2DCS. Such systems have been applied in Rb atoms and Cs2 molecules to induce electromagnetically-induced transparency 62, 63, and in 1,2-propanediol for enantiodetection 64. The introduction of these fields facilitates the generation of novel phenomena, allowing for the combination with other techniques or the execution of more refined operations. The RF formalism is used as a reference because it is a widely-used method that can provide analytical solutions with a clear physical picture. Although Lemmer et al. first applied the NHH method to 2DCS 55, they did not focus on the approach itself but rather on its application to study ion Coulomb crystals. In contrast, our work concentrates on the similarities and differences between these two methods in calculating 2DCS. We validate their consistency by comparing the 2DCS when they follow identical Liouville paths. We discuss their respective advantages and disadvantages, as well as their scopes of application. Particularly, we propose the quasi-Green function for interpreting the 2DCS by the NHH method, which significantly simplifies the analytical challenges. Then we use the NHH method to simulate the 2DCS of Rh(CO)2C5H7O2 (RDC) dissolved in hexane to further validate this approach. Overall, we believe that the NHH method provides a reliable description for 2DCS, particularly in driven or complex systems, and can overcome the limitations of the RF formalism in these scenarios.
@@
We first illustrate the RF formalism and the NHH method, in which the quasi-Green function is proposed as a counterpart to the Green function in the RF formalism. We then present their double-sided Feynman diagrams and the analytical expressions of the rephasing (RP) signal. We next provide a comparison of the Green and quasi-Green functions, the population dynamics, the conservation of the population and the 2DCS by the two approaches, focusing mainly on their differences in coherent and incoherent behavior. The advantages and disadvantages, as well as the difference and consistency of the two methods are discussed. We further apply the NHH method to RDC dissolved in hexane and discussed the results in relation to its energy levels and the existing experimental results. Finally, we summarize our results and offer a perspective on the application of the NHH method.
@@ 2.1 Response Function Formalism
We consider a three-level system including a ground state $|b\rangle$, an excited state $|e\rangle$, and a metastable state $|c\rangle$, as shown in Fig. 1(a). A control field with Rabi frequency $\Omega_{ec}$ is constantly applied to the system. When a probe field is introduced, the electromagnetically-induced transparency (EIT) effect is formed to suppress the dissipation due to quantum interference 65. For 2DCS, three incident pulses of the probe field induce the $|b\rangle$leftrightarrow $|e\rangle$ transition with Rabi frequency $\Omega_{be}$, and the signals to be measured are emitted from the system in a specific direction, as shown in Fig. 1(b). Between the energy levels $b$ and $e$, there exist bidirectional relaxations with a downhill rate $\Gamma_{1}$ and an uphill rate $\Gamma_{2}$. Basically, the Hamiltonian of the system without probe and control fields is
@@
$(j=b,e,c)$ where we assume $\hbar=1$ for simplicity.
Considering the external fields, $H_{\textrm{f}}(t)=H_{0}+H_{\textrm{int}}^{c}(t)$ is the total Hamiltonian including only the control field, and $H_{\textrm{p}}(t)=H_{0}+H_{\textrm{int}}^{c}(t)+H_{\textrm{int}}^{p}(t)$ is the total Hamiltonian including both the control field and the probe pulses. The interaction Hamiltonian $H_{\rm{int}}^{c}(t)$ is written as
$H_{\rm{int}}^{c}(t)=\Omega_{c}(t)|e\rangle\langle c|+{\rm h.c.},$ (2)
where $\Omega_{c}(t)=-\Omega_{ec}\exp(-i\nu_{c}t)/2$, with $\Omega_{ec}$ and $\nu_{c}$ being the Rabi frequency and the frequency of the control field respectively. On the other hand, the probe pulses are strong enough, as compared to the control field and the relaxations, to have the approximation $H_{\textrm{p}}(t)\simeq H_{0}+H_{\textrm{int}}^{p}(t)~{}(p=a,b,c)$. Meanwhile, the probe pulses are short enough to be considered as perturbations to the system. The Hamiltonian contributed by the probe pulses will be provided later in the NHH method when it is used.
Figure 1: (a) Schematic diagram of the three-level system. The energy levels $b$, $e$ and $c$ are the ground, the excited and the metastable states, respectively. The probe pulses denoted as the light-green line are applied to induce the $|b\rangle$leftrightarrow $|e\rangle$ transition with Rabi frequency $\Omega_{be}$, and the control field denoted as the dark-green line is resonantly coupled to the $|e\rangle$leftrightarrow $|c\rangle$ transition with Rabi frequency $\Omega_{ec}$. (b) Three probe pulses and the signal with their different directions. (c) Sequence of probe pulses (solid line), the control field (dotted line) and the signal (dashed line), in which $t_{c}~{}(c=1,2,3)$ is the time interval between two adjacent pulses.
The 2DCS are plotted using the third-order RF after double Fourier transformation. This function equals the third-order polarization $P^{(3)}(t)=\langle\mu\rho^{(3)}(t)\rangle$ in the semi-impulsive limit. Before considering the control field in $t_{c}$, the RF is expressed as 41, 66, 37, 67, 68
@@
$\langle\mu^{I}\left(t_{3}+t_{2}+t_{1}\right)\left[\mu^{I}\left(t_{2}+t_{1}\right),\left[\mu^{I}\left(t_{1}\right),\left[\mu^{I}\left(0\right),\rho^{I}\left(0\right)\right]\right]\right]\rangle,$
where $t_{1}$, $t_{2}$ and $t_{3}$ are the time intervals between two adjacent pulses, as shown in Fig. 1(c). $\rho^{I}(0)$ represents the initial equilibrium state of the system under the Hamiltonian $H_{0}$ in the interaction picture. Meanwhile, $\mu^{I}(t)=e^{iH_{0}t}\mu e^{-iH_{0}t}$ is the electric-dipole operator in the interaction picture, where $\mu=\mu_{be}|b\rangle\langle e|+\mu_{eb}|e\rangle\langle b|$ is the electric-dipole operator in the Schrödinger picture. By expanding the commutators, the third-order polarization is divided into eight terms with 41, 66, 37, 68
$\begin{split}R_{1}=&\langle\mu^{I}(t_{3}+t_{2}+t_{1})\mu^{I}(0)\rho^{I}(0)\mu^{I}(t_{1})\mu^{I}(t_{2}+t_{1})\rangle,\\ R_{2}=&\langle\mu^{I}(t_{3}+t_{2}+t_{1})\mu^{I}(t_{1})\rho^{I}(0)\mu^{I}(0)\mu^{I}(t_{2}+t_{1})\rangle,\\ R_{3}=&\langle\mu^{I}(t_{3}+t_{2}+t_{1})\mu^{I}(t_{2}+t_{1})\rho^{I}(0)\mu^{I}(0)\mu^{I}(t_{1})\rangle,\\ R_{4}=&\langle\mu^{I}(t_{3}+t_{2}+t_{1})\mu^{I}(t_{2}+t_{1})\mu^{I}(t_{1})\mu^{I}(0)\rho^{I}(0)\rangle.\end{split}$ (4)
and their hermitian conjugates. The incidence of three probe pulses from different directions allows for selective observation of certain combinations of the eight terms mentioned above. For example, the RP signal is emitted in the direction $\vec{k}_{s}=-\vec{k}_{a}+\vec{k}_{b}+\vec{k}_{c}$ as 41, 66, 37
@@
There are other methods to solve the signal starting from the RF, such as the generating function 67, 68, and here we present the approach using the Green function. $R_{n}$ $(n=1,2,3,4)$ and their hermitian conjugates, can be obtained by solving the quantum master equation. This equation describes quantum dynamics of the open system and is written as 69
@@
where $H$ is the total Hamiltonian, $A_{ij}=|i\rangle\langle j|$ is the quantum jump operator describing the process of jumping from $|j\rangle$ to $|i\rangle$, $\gamma_{j}^{\left(0\right)}$ is the pure-dephasing noise for each state, and $\mathfrak{L}\left(A_{ij}\right)\rho=\frac{1}{2}\left\{A_{ij}^{\dagger}A_{ij},\rho\right\}-A_{ij}\rho A_{ij}^{\dagger}$ with $\left\{A_{ij}^{\dagger}A_{ij},\rho\right\}=A_{ij}^{\dagger}A_{ij}\rho+\rho A_{ij}^{\dagger}A_{ij}$ being the anti-commutator. Assuming that the control field is in resonance with the $|e\rangle\leftrightarrow|c\rangle$ transition, the time-dependent term $e^{-i\nu_{c}t}$ in Eq. (2) is eliminated after the transformation to the rotating frame, i.e., the interaction picture. Thus, the time evolution of the elements of the density matrix in the interaction picture is as follows
$\begin{split}\dot{\rho}^{I}_{bb}=&\Gamma_{1}\rho^{I}_{ee}-\Gamma_{2}\rho^{I}_{bb},\\ \dot{\rho}^{I}_{ee}=&\frac{i}{2}\Omega_{ec}\left(\rho^{I}_{ce}-\rho^{I}_{ec}\right)-\Gamma_{1}\rho^{I}_{ee}+\Gamma_{2}\rho^{I}_{bb},\\ \dot{\rho}^{I}_{cc}=&-\frac{i}{2}\Omega_{ec}\left(\rho^{I}_{ce}-\rho^{I}_{ec}\right),\\ \dot{\rho}^{I}_{eb}=&\frac{i}{2}\Omega_{ec}\rho^{I}_{cb}-\gamma_{eb}\rho^{I}_{eb},\\ \dot{\rho}^{I}_{be}=&-\frac{i}{2}\Omega_{ec}\rho^{I}_{bc}-\gamma_{eb}\rho^{I}_{be},\\ \dot{\rho}^{I}_{ec}=&\frac{i}{2}\Omega_{ec}\left(\rho^{I}_{cc}-\rho^{I}_{ee}\right)-\gamma_{ec}^{+}\rho^{I}_{ec},\\ \dot{\rho}^{I}_{ce}=&-\frac{i}{2}\Omega_{ec}\left(\rho^{I}_{cc}-\rho^{I}_{ee}\right)-\gamma_{ec}^{+}\rho^{I}_{ce},\\ \dot{\rho}^{I}_{cb}=&\frac{i}{2}\Omega_{ec}\rho^{I}_{eb}-\gamma_{bc}\rho^{I}_{cb},\\ \dot{\rho}^{I}_{bc}=&-\frac{i}{2}\Omega_{ec}\rho^{I}_{be}-\gamma_{bc}\rho^{I}_{bc},\end{split}$ (7)
@@
When the non-diagonal elements $\rho_{ij}~{}(i\neq j)$ of the density matrix are treated perturbatively, and the Green functions are used to describe the time evolution of the density matrix elements, the RP signal is written in a more specific form 63
@@
where the Green function $\mathcal{G}_{ijkl}(t)=\rho^{I}_{ij}(t)$ for the initial condition $\rho^{I}_{mn}(0)=\delta_{mk}\delta_{nl}$. The double-sided Feynman diagrams for each term in Eq. (11) are shown in Fig. 2, with (a), (b), (c) and (d) corresponding to four terms in turn. The evolutions during $t_{2}$ with the initial and final states are explicitly demonstrated in the diagrams. The first and the last terms of Eq. (11) respectively correspond to $R_{2}$ and $R_{3}$ in Eq. (4). The other two terms, however, do not correspond to $R_{n}$ because they are induced by the population transfer during $t_{2}$. The evolution of the non-diagonal elements in $t_{1}$ and $t_{3}$ are described by the Green functions as
@@
while the evolution in $t_{2}$ is obtained by assuming two possible initial states after the first two pulses. In the situation where the system is at $|e\rangle$ at the beginning of the population time, and the non-diagonal elements of the density matrix are treated as perturbative, the Green functions for the system to remain at $|e\rangle$ or to evolve to $|b\rangle$ are respectively written as
@@
Eventually, by taking the real part of the double Fourier transformed RF, the RP signal and thus the 2DCS are obtained.
@@
We apply the NHH method to the same system in Fig. 1(a). Starting directly from the master Eq. (6), we drop the quantum-jump terms and approximately obtain
@@
where the NHH reads 70
@@
The NHH method is valid if two conditions are met. The first is that the dissipation and dephasing rates $\Gamma_{j}$ and $\gamma_{j}^{(0)}$ should be much smaller than the typical frequency of the Hamiltonian, while the other would be discussed later.
By the NHH method, the final state of the system in the interaction picture is written as 57, 64
$|\psi^{I}(t_{3},t_{2},t_{1})\rangle=U^{I}_{f}(t_{1}+t_{2}+t_{3}+\delta t_{a}+\delta t_{b}+\delta t_{c},$
@@
×$\times U^{I}_{p}\left(\delta t_{a},0\right)|\psi^{I}\left(0\right)\rangle,$ (22)
where $\delta t_{p}$ ($p=a,b,c$) is the time interval of the probe pulse, and $U^{I}_{p}$ ($U^{I}_{f}$) represents the evolution operator with (without) the probe pulses included in the interaction picture. As in the RF formalism, we consider that the probe pulses are strong enough relative to the control field, and their durations are short enough that both dissipation and dephasing during the excitations are neglected. As a result, the Hamiltonians during the probe pulses are explicitly written in the interaction picture as
$H^{I}_{p}=$ $-\frac{1}{2}\Omega_{be}e^{-i\vec{k}_{p}\cdot\vec{r}}|b\rangle\langle e|+{\rm h.c.}$ (23)
Taking dissipation and dephasing into account, the Hamiltonians during the free evolutions read
@@
The Rabi frequency $\Omega_{ij}$ is given by $-\vec{d}_{ij}\cdot\vec{E}$, where $\vec{E}$ is the electric field of the laser and $\vec{d}_{ij}$ is the transition dipole moment between state $i$ and $j$. Thus, the selectivity through laser polarization is accounted for and directly reflected in $\Omega_{ij}$.
Hereinto, we approximate the probe pulses as square pulses. Under $H^{I}_{f}$, the time evolutions $U_{f}^{I}(t_{c}+t_{0},t_{0})=e^{iH_{0}(t_{c}+t_{0})}U_{f}^{S}(t_{c}+t_{0},t_{0})e^{-iH_{0}t_{0}}=e^{iH_{0}t_{c}}U_{f_{I}}(t_{c})$, where $t_{0}$ represents the starting time of the free evolution, and $U_{f}^{S}$ is the time evolution operator in the Schrödinger picture. Then $U_{f_{I}}(t_{c})$ for initial states $|b\rangle$, $|e\rangle$ and $|c\rangle$ are respectively written as
$U_{f_{I}}(t_{c})|b\rangle$ $=C_{bb}(t_{c})|b\rangle,$ (25)
$U_{f_{I}}(t_{c})|e\rangle$ $=C_{ee}(t_{c})|e\rangle+C_{ce}(t_{c})|c\rangle,$ (26)
$U_{f_{I}}(t_{c})|c\rangle$ $=C_{ec}(t_{c})|e\rangle+C_{cc}(t_{c})|c\rangle,$ (27)
and
@@
Meanwhile, considering $\delta t_{p}\simeq 0$, the time evolutions $U_{p}^{I}(\delta t_{p}+t_{0},t_{0})=U_{p_{I}}(\delta t_{p})$ with $t_{0}$ denoting the starting time of the probe pulses. The $U_{p_{I}}(\delta t_{p})$ under $H^{I}_{p}$ for three initial states are then respectively written as
$U_{p_{I}}(\delta t_{p})|b\rangle=$ $N_{p}\left(|b\rangle+\beta_{p}e^{i\vec{k_{p}}\vec{r}}|e\rangle\right),$ (36)
$U_{p_{I}}(\delta t_{p})|e\rangle=$ $N_{p}\left(\beta_{p}e^{-i\vec{k_{p}}\vec{r}}|b\rangle+|e\rangle\right),$ (37)
$U_{p_{I}}(\delta t_{p})|c\rangle=$ $|c\rangle,$ (38)